# Structural Change, Employment, and Inequality in Europe: an Economic Complexity Approach

### v1

@@
Bernardo Caldarola Corresponding author. Email: caldarola@merit.unu.edu Enrico Fermi Research Center, Rome, Italy UNU-MERIT, Maastricht, the Netherlands School of Business and Economics, Maastricht University, Maastricht, the Netherlands Dario Mazzilli Enrico Fermi Research Center, Rome, Italy Aurelio Patelli Enrico Fermi Research Center, Rome, Italy Angelica Sbardella Enrico Fermi Research Center, Rome, Italy
Structural change consists of industrial diversification towards more productive, knowledge-intensive activities. However, changes in the productive structure bear inherent links with job creation and income distribution. In this paper, we investigate the consequences of structural change – defined in terms of labour shifts towards more complex industries – on employment growth, wage inequality and functional distribution of income. The analysis is conducted for European countries using data on disaggregated industrial employment shares over the period 2010 – 2018. First, we identify patterns of industrial specialisation by validating a country-industry industrial employment matrix using a bipartite weighted configuration model (BiWCM). Secondly, we introduce a country-level measure of labour-weighted Fitness, which can be decomposed in such a way as to isolate a component that identifies the movement of labour towards more complex industries – the structural change component. Thirdly, we link structural change to i) employment growth, ii) wage inequality, and iii) labour share of the economy. The results indicate that our structural change measure is associated negatively with employment growth. However, it is also associated with lower income inequality. As countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear. Finally, structural change predicts a higher labour ratio of the economy; however, this is likely to be due to the increase in salaries rather than by job creation.
Keywords: Structural Change, Economic Complexity, Inequality, Employment JEL Codes: D63, E24, J31, O11, O52
@@ 1 Introduction
Structural change can be defined by the reallocation of economic activity across bread sectors of the economy (Lewis 1954, Syrquin 1988). The shift of the relative importance of industrial employment or output towards increasingly more productive activities that defines the process of structural change is also an engine of long-term, sustained economic growth (Herrendorf et al. 2014, 2022). However, changes in the productive structure are also inherently associated with the distribution of income, as highlighted in the seminal work by Kuznets (1955). For instance, when countries upgrade their productive structure, they also move to sectors with higher knowledge intensity, and lower labour requirements. This may have consequences on the wage and functional distribution of income, as well on the creation of new jobs and the destruction of old ones.
In this paper, we use the case of Europe to study the relationship between the movement of labour to increasingly knowledge intensive activities – defined here as structural change – with employment growth, wage inequalities, and the functional distribution of income between labour and capital.
As a matter of fact, in the last decades within country inequality has been growing in Europe and in high-income countries. At the same time, these countries have de-industrialised during the last wave of ICT-led globalisation, leading to the substitution of manufacturing with jobs in complex, knowledge intensive industries – such as business services – contributing to increasing inequalities (Antonelli & Tubiana 2020). The mechanism underlying this trend has been explored further in the empirical literature. The Skill-Biased Technological Change (SBTC) literature (Autor et al. 2003, Autor 2022, Acemoglu & Restrepo 2022) has examined the impact of the main driver of structural change – technological change – on employment and distribution of earnings, suggesting that waves of technological change tend to complement only more educated workers in knowledge intensive industries, leaving less educated workers behind.
@@
Moreover, structural change bears a strong linkage with employment creation. As economies transform in their productive structures, they create winners and losers (Schumpeter 1934). The current wave of technological change may not reduce total employment (Dauth et al. 2021), but induces labour market adjustments at the disadvantage of low-skilled workers, whose share over total workers is reduced (Graetz & Michaels 2018). In the case of other technologies, such as ICT, these may have been the culprits of jobless growth (Brynjolfsson & McAfee 2011, Frey & Osborne 2017) and contributed to the shrinking of the number of middle-skill jobs (Goos et al. 2009, 2014).
@@
In order to identify the degree of knowledge intensity of European industries we employ an Economic Complexity (EC) perspective. Complexity metrics aim at quantifying the capability structure of countries by looking at their productive basket, preserving information on their specialisation and diversification patterns rather than by simply aggregating their economic output (Hausmann & Hidalgo 2009, Tacchella et al. 2012). By the same token, EC measures allow to identify the capability (knowledge) intensity (Lo Turco & Maggioni 2022) of individual products, industries, or technologies by considering their ubiquity – a proxy for the diffusion of the capabilities that such activities require – weighted by the capability endowment of the countries specialised in such activities – i.e., the countries’ Fitness. Given the strong link of EC metrics with economic diversification, and in particular of the Economic Fitness and Complexity (EFC) approach (Tacchella et al. 2012), they have proved to be a useful tool to describe a long-term dynamic such a structural change (Freire 2021, Castañeda et al. 2022, McNerney et al. 2023), being also highly predictive of GDP growth (Tacchella et al. 2018). Following the analogy between structural change and complexity, the latter can be used to explain the inequality trends that result of changes in the sectoral composition of the economy. The literature on complexity and inequalities (Sbardella et al. 2017, Hartmann et al. 2017, Hartmann & Pinheiro 2022) indicates that higher complexity of the productive structure is associated to lower inequalities at the country level. It is also well known that knowledge concentrates in space (Balland et al. 2020, 2022), turning geography into a useful lens through income distribution can be analysed. In fact, at the regional level the sign of this relationship flips (Sbardella et al. 2017), as only regions with the strongest knowledge basis are able to diversify into highly complex industries, leaving laggards behind (Pinheiro et al. 2022). Looking at the relationship between complexity and employment, Adam et al. (2023) finds that higher complexity is linked to job creation. Where existing jobs are made redundant, it has been shown that re-employment is easier in presence of related industrial variety at the regional level (Hane-Weijman et al. 2018).
One important feature of EC methods is that these allow to quantify the capability requirements of finely disaggregated industries, offering a granular description of the process of structural change both across and within broad sectors of the economy. In order to offer an encompassing assessment of the complexity of European industries and of the labour flows across them, we depart by traditional approach of the EC literature – which infers the productive structure of countries using export data – by using data on industrial employment in European countries between 2010 and 2019. First, we consider the use of employment data to measure industrial Complexity and countries’ employment Fitness as this allows to include also non-tradeable activities, most of which are in the service sector. This is particularly relevant considering that European economies have heavily shifted their productive structure towards the tertiary industries along with the process of globalisation. Secondly, the focus on employment represents a crucial element in the conceptualisation of structural change as the change in relative shares of the economy. Using employment data to measure Economic Fitness and Complexity respectively at the country and industry level resonates with this conceptual framework, and allows to outline the structure of capabilities of labour as a productive input.
@@
In the present study, the analysis of the relationship between structural change, inequality, and employment creation is conducted for European countries over the period 2010 -– 2018. The empirical approach relies on information on European industrial employment provided by Eurostat’s Structural Business Survey data, measured at a fine level of disaggregation (NACE 4 digits) across manufacturing and service industries. This information is matched with data on the distribution of wages at the country level, provided by the ILO, and with data on the labour share of the economy and employment growth, obtained by ARDECO. Using this newly constructed dataset, we start by identifying patterns of industrial specialisation by validating the country-industry industrial employment matrix using a bipartite weighted configuration model (BiWCM), overcoming some limitations imposed by the Balassa (1965) method. Secondly, we introduce a measure of labour-weighted Fitness, which sums up complexity of industries weighted by their employment share. This measure can be decomposed in such a way as to identify the contribution to changes in labour-weighted Fitness coming from the movement of labour towards more complex industries. We identify such component as the one linked to structural change from a labour viewpoint – the structural change component. Thirdly, we link the structural change component to a number of economic outcomes at the country level: i) employment growth, ii) wage inequality, and iii) functional distribution of income (labour share of the economy). In this way, we analyse the distributional consequences of structural transformation – framed as the movement of labour to more complex industries – and its job-creating potential.
@@
An analysis of the specialisation patterns in European countries, conducted by inspecting the significant links of the country-industry specialisation matrix generated by the BiWCM, reveals that the most diversified countries have progressively moved away from specialisation in low-complexity, labour-intensive industries. Moreover, our empirical analysis explores the association between the structural change component of our labour-weighted Fitness measure -– derived from the movement of labour towards more complex industries -– to employment growth, wage inequality, and the functional distribution of income (labour share of the economy). The preliminary results of OLS panel regressions with country and time fixed effects indicate that our structural change measure is associated negatively with employment growth, corroborating the evidence that highly complex industries have lower labour requirements. However, it is also associated with lower income inequality – measured in terms of the ratio of average wages in the ninth and first deciles of the wage distribution. As countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear, making the first decile of the salary distribution go up. Finally, structural change predicts a higher labour ratio of the economy; however, this is likely to be due to the increase in salaries in highly complex industries rather than by job creation. Very interestingly, the structural change component is the only one to be significant in the regressions, while the increase in within-sector complexity, and the labour-weighted Fitness of countries do not explain our outcome variables. Results are robust to a battery of robustness checks, including leave-one-out regressions and a specification where the structural change component of the labour-weighted Fitness is replaced by a similar component obtained decomposing a labour-weighted entropy measure.
@@
The remainder of this paper is organised as follows. In Section 2 we delve into the Economic Fitness and Complexity framework, and we introduce a new measure – the Labour-Weighted Fitness (LWF) the variation of which can be decomposed to extract a structural change component. In Section 3 we outlined the data sources used in the analysis, which in turn is described in the following Section (4). Section 5 illustrates the preliminary results of the analysis, and Section 6 concludes.
2 Analytical Framework
@@
In this section, we will outline the analytical framework used to derive a measure of structural change based on the reallocation of labour to complex industries. First, we will define countries’ Fitness and industrial complexity according to the EFC framework (Tacchella et al. 2012). Second, we introduce a new measure – the Labour Weighted Fitness – from which we derive a structural change component.
2.1 Structural change, Fitness and Complexity
Economic Complexity is a framework that builds on earlier evolutionary and institutional literature (Cimoli & Dosi 1995, Hirschman 1958, Teece et al. 1994) to tackle the complexity of economic systems. It describes the economy as an evolutionary process of globally interconnected ecosystems. The main recent advance with respect to the earlier literature is the use of newly developed network science and other methods to investigate complex and dynamical systems (Hausmann & Klinger 2006, Hausmann & Hidalgo 2009, Tacchella et al. 2012) to separate the random noise from the underlying signal. The Economic Complexity framework shifts the focus from aggregate quantities (What is the GDP of the country?) to a more disaggregated view (In which industrial sectors do countries specialise?), with the aim of providing information that is complementary to more traditional analysis. This shift creates the opportunity to create a consistent framework that allows several cross-cutting themes to be addressed and provides a quantitative answer to several policy relevant questions that could otherwise only be answered qualitatively or by ad hoc metrics.
In particular, EC methods have contributed “to the study of structural change by addressing questions related to how the diversification of countries and the patterns of trade specialization affect and are affected by economic growth, and how diversification and trade specialization change over time, and the determinants of such change” (Freire 2021, p 214), providing useful tools to study disaggregated aspects of structural change.

### v2

@@
Bernardo Caldarola Corresponding author. Email: bernardo.caldarola@ec.europa.eu European Commission, Joint Research Center, Seville, Spain Enrico Fermi Research Center, Rome, Italy UNU-MERIT and Maastricht University, Maastricht, the Netherlands Dario Mazzilli European Commission, Joint Research Center, Seville, Spain Aurelio Patelli European Commission, Joint Research Center, Seville, Spain Angelica Sbardella European Commission, Joint Research Center, Seville, Spain
Structural change consists of industrial diversification towards more productive, capability-intensive activities, bearing inherent links with job creation and income distribution. This paper investigates the consequences of structural change –- defined in terms of labour shifts towards more complex industries –- on employment growth, wage inequality, and the functional distribution of income between labour and capital. The analysis is conducted for European countries using data on granular sector-level employment shares over the period 2010-2018. Structural change is described using a country-level measure that captures the movement of workers towards more complex industries. The findings indicate that the complexity-based measure of structural change is associated negatively with employment growth. However, it is also associated with lower income inequality: as countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear. Finally, structural change predicts a higher labour ratio of the economy; however, this is likely to be due to movement of the workforce into higher-wage, more complex sectors rather than to job creation.
Keywords: Structural Change, Economic Complexity, Wage Inequality, Labour Share, Employment Growth JEL Codes: D63, E24, J31, O11, O52
@@ 1 Introduction
Structural change describes the process of reallocation of economic activity across economic sectors (Lewis, 1954; Syrquin, 1988; Matsuyama, 2008) with heterogeneous levels of productivity, technological content, and labour requirements (Baumol, 1967) which drives long-term economic growth (Herrendorf et al., 2014, 2022). Commonly, structural change is quantified by changes in the relative importance of industries – measured by their shares in total employment or output. This process has been shown to rest on the underlying accumulation of increasingly sophisticated productive capabilities (Verspagen, 1991; Abramovitz, 1995; Cimoli et al., 2009), which allows economies to innovate and diversify through the recombination of existing knowledge (Hausmann and Hidalgo, 2009; Tacchella et al., 2012; Rodrik and others, 2013; van Dam and Frenken, 2022; Inoua, 2023; Teixeira and Pinto, 2026).
However, structural change is neither a uniform nor an automatic process. Different varieties of structural transformation may lead to distinct outcomes depending on the sectors into which economies diversify and the institutional contexts in which these transformations unfold, according to which its gains are distributed across workers and between labour and capital. This is particularly relevant in advanced economies, where structural change has increasingly taken the form of deindustrialisation, tertiarisation, and the expansion of knowledge-intensive sectors rather than the classical shift from agriculture to manufacturing (Galbraith, 2011; Herrendorf et al., 2014; Rodrik, 2016). Against this backdrop, this paper focuses on one specific aspect of this broader process: how structural change, conceptualised as the reallocation of labour towards more complex and capability-intensive sectors, is associated with employment creation and income distribution in European economies.
In his seminal work, Kuznets (1955) argued that inequality would initially rise and then decline as economies transition from traditional to manufacturing sectors. In this framework, income disparities arise from productivity differentials across sectors: as long as low and high productivity sectors coexist, aggregate inequality reflects the income gap between them. Inequality would initially rise as a modern sectors start to grow, but as sectoral reallocation converges towards the higher-productivity sectors, inequality is predicted to decline. Although this inverted-U shape relationship between development and inequality proposed by Kuznets has been widely debated (Atkinson and Brandolini, 2006; Hellier and Lambrecht, 2013; Galbraith et al., 2025), the broader insight that sectoral reallocation and income distribution are structurally linked remains influential. The multiplicity of definitions of both “structural change” and “inequality”, together with the diversity of empirical patterns observed across countries and historical periods, has prevented the emergence of a unified consensus regarding the direction and magnitude of this relationship (Cosentino De la Vega, 2023; Förster and Tóth, 2015; Kanbur, 2000). Moreover, several mechanisms are at play along with sectoral changes, with an effect of income distribution and employment creation. These include globalisation, technological change, labour market institutions, political processes, tax regimes, and demographics (Förster and Tóth, 2015). Nevertheless, the extant literature on this topic has not built a direct link between the accumulation of (increasingly rare) capabilities as sectoral changes unfold and the distribution of income.
In this paper, we contribute to this debate by investigating the labour consequences of knowledge recombination towards more sophisticated capabilities. To this aim, we use a definition of structural change that moves beyond measures based on productivity or sectoral shares, and builds explicitly on capabilities accumulation a fundamental notion of structural change. Our definition and empirical operationalisation of structural change pins down the movement of labour to industries that require more sophisticated capabilities using the Economic Complexity framework (Hausmann and Hidalgo, 2009; Tacchella et al., 2012). First, we construct an employment-based measure of Economic Fitness which quantifies the diversification of industries specialisation of a country, weighted by the sophistication of such industries. Second, we separate changes in labour-weighted Fitness between i) the reallocation of labour towards more sophisticated and capability-intensive sectors, and ii) changes in industry sophistication due to changes in the capabilities required by the industry itself. We relate the former – the structural change component – to employment growth, wage inequalities, and the functional distribution of income between labour and capital. The proposed approach offers a perspective on structural change and inequalities that links directly the process of capability accumulation that underlies structural transformations to the patterns in employment creation and labour income distribution associated to it.
Understanding these dynamics is particularly relevant in light of the profound transformations observed in European and other high-income economies over recent decades, which have been accompanied by rising within-country inequality, declining employment growth, and a falling labour share of income. Within-country inequality has been on the rise in Europe and in other high-income countries (Piketty, 2014; Piketty et al., 2018), with increasing between-industry wage dispersion (Haltiwanger et al., 2024). Increasing wage inequalities have been accompanied by a decline in the labour share of income across European countries and the US alike (Jorgenson and Timmer, 2011; Karabarbounis and Neiman, 2014; Stockhammer, 2017; Karabarbounis, 2024), redistributing wealth away from workers towards capital owners. At the same time, most high-income countries have de-industrialised during the last wave of ICT-led globalisation, leading to the substitution of manufacturing with jobs in complex, knowledge-intensive industries, and contributing to increasing inequalities (Cirillo, 2018; Antonelli and Tubiana, 2020).
The mechanisms underlying these trends has been widely explored in the empirical literature. The Skill-Biased Technological Change (SBTC) literature (Autor et al., 2003; Autor, 2022; Acemoglu and Restrepo, 2022) has examined the impact of technological change, the main driver of structural change, on employment and on the distribution of earnings, suggesting that recent waves of technological change, especially ICT capital goods, tend to complement high-skilled workers in knowledge-intensive industries, leaving low-skilled workers behind. Later this literature has argued that digital technology-driven transformations may affect not only the relative demand for skilled labour, but also the occupational structure itself, giving rise to increasing job polarisation, that is, to the disappearance of middle-wage jobs (Autor et al., 2006; Goos et al., 2009) which are more easily replaced by capital (Acemoglu and Loebbing, 2026). Additionally, offshoring, the expansion of global value chains, and import competition from lower-wage countries have been seen as contributing to reducing the demand for low- and middle-skilled labour in advanced economies, thereby reinforcing wage inequality and labour market polarisation (Feenstra and Hanson, 1996; Autor et al., 2013; Timmer et al., 2014). Alongside these market-based explanations, the institutional literature interprets the distributive outcomes associated with structural change not as the mechanical result of technology and trade alone, but also of political and institutional changes that shape wage-setting mechanisms and the division of income between labour and capital. In this perspective, rising inequality and the decline in the labour share reflect not only shifts in labour demand, but also the erosion of workers’ bargaining power through declining unionisation, labour market deregulation, weaker labour standards, and welfare state retrenchment (Mishel and Bivens, 2021; Stansbury and Summers, 2020; Stockhammer, 2012).
@@
Another crucial feature of structural change is its linkage with employment creation. As economies transform in their productive structures, new jobs are created while some others are made redundant (Schumpeter, 1934). As the employment effects of structural change depend on the labour intensity of expanding sectors, structural upgrading may be associated with lower employment growth even in the presence of rising productivity (Kucera and Roncolato, 2012). On the one hand, the current wave of automation may not reduce total employment (Dauth et al., 2021), but may induce labour market adjustments at the disadvantage of low-skilled workers, whose share over total workers is reduced (Graetz and Michaels, 2018). On the other hand, ICTs may have been among the culprits of jobless growth (Brynjolfsson and McAfee, 2011; Frey and Osborne, 2017) and contributed to the shrinking of the number of middle-skill jobs (Goos et al., 2009, 2014) due to the lower labour requirements of ICT intensive industries.
@@
In order to identify variety and uniqueness of capabilities required to achieve industrial upgrading we employ an Economic Complexity (EC) perspective. Economic Complexity metrics aim at quantifying the hierarchical capability structure of countries by looking at their productive basket, preserving information on their specialisation and diversification patterns rather than by simply aggregating their economic output (Hausmann and Hidalgo, 2009; Tacchella et al., 2012; Diodato et al., 2024). By the same token, EC measures allow to identify the capability requirements (Lo Turco and Maggioni, 2022) of individual products, industries, or technologies. Given the strong link of EC metrics with economic diversification, and in particular of the Economic Fitness and Complexity (EFC) approach (Tacchella et al., 2012), they have proved to be a useful tool to describe a long-term dynamic such a structural change (Freire, 2021; Castañeda et al., 2022; McNerney et al., 2025), being also highly predictive of GDP growth (Tacchella et al., 2018).
Following the analogy between structural change and complexity, the latter can be used to explain the inequality trends that result from changes in the sectoral composition of the economy. The literature on complexity and inequalities (Sbardella et al., 2017; Hartmann et al., 2017; Hartmann and Pinheiro, 2022) indicates that more complex activities are better remunerated, and that higher complexity of the productive structure is associated to lower inequalities at the country level (Sbardella et al., 2017; Hartmann et al., 2017; Hartmann and Pinheiro, 2024; Aufiero et al., 2024). However, this relationship may present non-linearites and also reflect cross-country differences in institutional settings, including the inclusiveness of labour market and welfare institutions (Hartmann et al., 2017; Sbardella et al., 2017; Napolitano et al., 2022; Hartmann and Pinheiro, 2024). Looking at the relationship between complexity and employment, Adam et al. (2023) finds that higher complexity is linked to job creation. Where existing jobs are made redundant, it has been shown that re-employment is easier in presence of related industrial variety at the regional level (Hane-Weijman et al., 2018).
@@
We measure the relative importance of finely disaggregated industries (NACE 4 digits) using data on industrial employment in European countries between 2010 and 2018111Similar efforts to compute Economic Complexity indices for countries or regions have been performed, among others, by Sbardella et al. (2017); Caldarola (2022); Pinheiro et al. (2022) obtained from Eurostat’s Structural Business Survey data and compute a measure of Economic Fitness weighted by labour shares. Using data on industrial employment allows us to include also non-tradeable activities, most of which are in the service sector. [^2] Mishra et al. (2020) have also been able to include services activities in the computation of Economic Fitness, however limited to tradable ones. This is particularly relevant considering that European economies have heavily shifted their productive structure towards the tertiary industries along with the process of globalisation. Secondly, the focus on employment represents a crucial element in the conceptualisation of structural change as the change in relative shares of the economy. Data on the individual and functional distribution of income are obtained, respectively, from ILO and ARDECO; the latter is also used to measure employment growth in the observed period. First, we identify patterns of industrial specialisation by validating the country-industry industrial employment matrix using a null model approach – the Bipartite Weighted Configuration Model (Bruno et al., 2023) – that overcomes some limitations imposed by Balassa’s (1965) Revealed Comparative Advantage index. Secondly, we introduce a measure of labour-weighted Fitness, which sums up the complexity of industries weighted by their employment share. This measure can be decomposed in such a way as to identify the contribution to changes in labour-weighted Fitness coming from the movement of labour towards more complex industries. We identify such component as the one linked to structural change from a labour viewpoint –- the structural change component. Thirdly, we link the structural change component to a number of economic outcomes at the country level: i) employment growth, ii) wage inequality, and iii) the functional distribution of income (labour share of the economy). In this way, we analyse the distributional consequences of structural transformation – framed as the movement of labour to more complex industries – and its job-creating potential.
@@
Our analysis of the specialisation patterns in European countries reveals that the most diversified countries have progressively moved away from specialisation in low-complexity, labour-intensive industries. The results of a fixed effect panel regressions indicate that our structural change measure is negatively associated with employment growth, corroborating the evidence that highly complex industries have lower labour requirements. However, it is also associated with lower income inequality –- measured in terms of the ratio of average wages in the ninth and first deciles of the wage distribution. As countries move to more complex industries, they drop the least complex ones, so the (low-paid) jobs in the least complex sectors disappear, making the first decile of the salary distribution go up. Finally, structural change is associated with a higher labour share of the economy, likely mirroring a reallocation of labour towards complex and higher-paying industries rather than a employment expansion. Interestingly, the structural change component is the only one to be significant in the regressions, while the increase in within-sector complexity, and the labour-weighted Fitness of countries do not explain our outcome variables. Results are robust to a battery of robustness checks, including leave-one-out regressions and a specification where the structural change component of the labour-weighted Fitness is replaced by a similar component obtained decomposing a labour-weighted entropy measure, that takes into account the breath of sectoral specialisation but not its complexity. These findings imply the need for industrial, innovation, and labour-market policies capable of supporting transitions towards more complex activities while preserving job creation capacity and ensuring that the gains from structural transformation are distributed in an inclusive way.
@@
The remainder of this paper is organised as follows. In Section 2 we delve into the Economic Fitness and Complexity framework, and we introduce a new measure, the Labour-Weighted Fitness (LWF) the variation of which can be decomposed to extract a structural change component. In Section 3 we outline the data sources used in the analysis, which in turn is described in the following Section (4). Section 5 illustrates the main results of the analysis, and Section 6 concludes.
2 Structural change and Economic Complexity
@@
In this section, we outline the analytical framework used to derive a measure of structural change based on the reallocation of labour to complex industries. First, we will define country Fitness and industry complexity according to the EFC framework. Second, we introduce a new measure – the Labour Weighted Fitness – from which we derive the structural change component of interest.
2.1 The Economic Complexity framework
The Economic Complexity framework builds on earlier evolutionary and institutional traditions that emphasise the role of productive capabilities, sectoral composition, and cumulative learning in shaping development trajectories (Cimoli and Dosi, 1995; Hirschman, 1958; Teece et al., 1994). At its core lies the idea that the production of goods and services requires combinations of highly specific and complementary underlying capabilities (Abramovitz, 1995; Hausmann and Rodrik, 2003), such as skills, technologies, organisational routines and institutions. Because these capabilities cannot be measured directly, the Economic Complexity approach infers their distribution from observable patterns of specialisation across economic activities. By examining the industries or products in which countries display a comparative advantage, it becomes possible to recover information about the underlying structure of capabilities embedded in their economies (Hausmann and Klinger, 2006; Hausmann and Hidalgo, 2009; Tacchella et al., 2012). This perspective shifts the focus from aggregate outcomes, such as GDP, to the composition of productive activity, asking in which industries do countries specialise and what does this reveal about their underlying capabilities. To do so, Economic Complexity methods represent specialisation profiles as networks linking countries to the activities in which they display a comparative advantage, and then exploits the structural information embedded in these networks to construct ad hoc metrics grounded in complex systems theory. By leveraging highly disaggregated data and tools from network science and complex systems analysis, the Economic Complexity framework provides a coherent empirical approach to study diversification, technological upgrading, and structural transformation, complementing more traditional (and aggregate) analyses. In particular, EC methods have contributed “to the study of structural change by addressing questions related to how the diversification of countries and the patterns of trade specialization affect and are affected by economic growth, and how diversification and trade specialization change over time, and the determinants of such change” (Freire, 2021, p 214).